40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 245 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  4775   Tue May 31 17:30:42 2011 steveUpdateSUSETMY damping restored

ETMY's watch dogs were found tripped. They were restored.

  4774   Tue May 31 16:07:57 2011 Larisa ThorneConfigurationElectronicsSeismometer Box Update

 (Continuation of this)

 

I plugged the circuit into the LISO program to generate the graphs below....the first graph is a plot of frequency (f, in Hz) versus gain (in dB), and frequency (f, Hz) versus phase (in degrees). Also included is the second graph, which is a noise plot of all circuit parts which contribute to the total noise of the circuit.

 

The only issue I had was that two of the op amps I'd picked (see third attachment for the original circuit diagram) for the circuit were not in LISO's op amp library. So I replaced THS4131 (from the voltage buffer part) and AD826 (from the ADC driver part) with AD797 and LT1037, respectively in order to generate the plots below....   

 

There are notes calling the AD797 "ultra low noise, low distortion", whose data sheet can be found here: AD797 

Notes also call LT1037 "low noise, high speed precision op amp", whose data sheet can be found here: LT1037

 

I've put these in temporarily only, as I don't know if they are appropriate choices for the job or even if we have them. Suggestions?

Attachment 1: SeisBoxLISOplot1.pdf
SeisBoxLISOplot1.pdf
Attachment 2: SeisBoxLISOplot2.pdf
SeisBoxLISOplot2.pdf
Attachment 3: STS2diagram_original.pdf
STS2diagram_original.pdf
  4773   Tue May 31 15:45:37 2011 JamieUpdateCDSc1iscey IOchassis powered off for some reason. repowered.

We found that both of the c1iscey models (c1x05 and c1scy) were unresponsive, and weren't coming back up even after reboot.  We then found that the c1iscey IOchassis was actually powered off.  Steve's accepts some sort of responsibility, since he was monkeying around down there for some reason.  After powerup and reboot, everything is running again.

  4772   Tue May 31 14:29:00 2011 jzweizigUpdateCDSframes

There seems to be something strange going on with the 40m frame builder.
Specifically, there is a gap in the frames in /frames/full near the start of
each 100k second subdirectory. For example, frames for the following times are missing:

990200042-990200669
990300045-990300492
990400044-990400800
990500032-990500635
990600044-990600725
990700037-990700704
990800032-990800677
990900037-990900719


To summarize, after writing the first two frames in a data directory, the next ~10 minutes of frames are usually missing. To make matters worse (for
the nds2 frame finder, at least) the first frame after the gap (and all successive frames) start at an arbitrary time, usually not aligned to a 16-second boundary. Is there something about the change of directories that is causing the frame builder to crash? Or is the platform/cache disk too slow to complete the directory switch-over without loss of data?

  4771   Tue May 31 11:34:13 2011 steveSummarySAFETYsafety glasses checked

  1064 nm transmison were measured of 40m safety glasses as shown . Their performance did not degrade. They are as good as their labels.

Attachment 1: P1070823.JPG
P1070823.JPG
  4770   Tue May 31 11:26:29 2011 josephbUpdateCDSCDS Maintenance

1) Checked in the changes I had made to the c1mcp.mdl model just before leaving for Elba.

2) The c1x01 and c1scx kernel modules had stopped running due to an ADC timeout. 

According to dmesg on c1iscex, they died at 3426838 seconds after starting (which corresponds to ~39 days).  "uptime" indicates c1iscex was up for 46 days, 23 hours. So my guess is about 8 days ago (last Monday or Tuesday),  they both died when the ADCs failed to respond quick enough for an unknown reason.

I used the kill scripts (in /opt/rtcds/caltech/c1/scripts/) to kill c1spx, c1scx, and c1x01.  I then used the start scripts to start c1x01, then c1scx, and then finally c1spx.  They all came up fine.

Status screen is now all green.  I renabled damping on ETMX and it seems to be happy. A small kick of the optic shows the approriately damped response.

  4769   Mon May 30 23:14:27 2011 valeraUpdateASCY arm initial alignment

I closed all 8 dither loops for the Y arm initial alignment: 2x2 centering servo (this worked before) and 2x2 input beam servo for both pitch and yaw.

So far it looks pretty good - the error points go to zero and the arm power goes up to 1.

The offloading to the alignment biases and the PZTs is not yet automated.

Today the PMC, MC, and Y arm were very cooperative and a pleasure to work with.

  4768   Fri May 27 17:52:53 2011 steveUpdateLSCLSC rack cables strain relieved & labeled

LSC rack 1Y2 cables are strain relieved and labeled. Spare and/or obsolete cables are laid out on the top of the beam tube and on the outside of the rack.

The POY 110 MHZ demodboard has a very touchy position in the VME crate. Watch out for it! It has to be fixed.

  4767   Thu May 26 17:10:21 2011 JamieConfigurationLSCNew digital lockin added to LSC model/screen

Quote:

A future improvement would be to modify the DCPD input matrix to be able to route those signals to the lockin as well.  This is actually currently possible since the DCPD input matrix is just a subset of the full input matrix, but it's not available via medm yet.

 I went ahead and added the lockin output to the DCPD input matrix.

 

  4766   Wed May 25 20:12:55 2011 JamieConfigurationLSCNew digital lockin added to LSC model/screen

Quote:

This is OK....but, the input matrix should come from the same place as the regular input matrix: i.e. it should be just another row like CARM, DARM, etc. rather than have its own screen.

 You're absolutely right.  That was a brain-fart oversight.  I fixed the model so that the input from the lockin comes from another output row in the RFPD input matrix.  I then fixed the C1LSC medm screen accordingly:

lsc-lockin-adl.png

This is obviously much simpler and more straight-forward.

A future improvement would be to modify the DCPD input matrix to be able to route those signals to the lockin as well.  This is actually currently possible since the DCPD input matrix is just a subset of the full input matrix, but it's not available via medm yet.

  4765   Wed May 25 19:19:11 2011 JamieConfigurationCDS!!!CHECK IN YOUR MODELS!!!

!!!CHECK IN MODEL CHANGES!!!

Today I found three models that were modified, but not checked in to the SVN repository:

M       sus/c1/models/lib/sus_single.mdl
M       isc/c1/models/c1lsc.mdl
M       isc/c1/models/c1mcp.mdl

I checked in the c1lsc model, since I think it was just the change that Kiwamu made in http://nodus.ligo.caltech.edu:8080/40m/4749.  I left the others, since I have no idea what they are or who made them.

Please please please remember to commit your model changes to the SVN after you're done.  This is particularly important for important models, such as c1lsc.  If you don't check in your changes I can pretty much guarantee that at some point you will loose them.

 

  4764   Wed May 25 19:03:59 2011 JamieConfigurationCDSUpdate rtcds checkout of cds_user_apps with new top-level directory names.

The top-level subsystem subdirectories in the cds_user_apps source repository were renamed today to be all lower case.  This required checking out the new directory and updating all of the model links in /opt/rtcds/caltech/c1.  Here is how I updated the cds_user_apps working tree:

cd /opt/rtcds/caltech/c1/userapps
mv trunk{,.bak}
svn update --depth=files trunk
svn update --depth=empty trunk/{cds,isi,isc,psl,sus}

I then fixed the links in the /opt/rtcds/caltech/c1/core/release/src/epics/simLink directory:

for link in $(find . -maxdepth 1 -type l); do ln -sf $(readlink $link | tr [:upper:] [:lower:]) ; done

A couple of things had to be cleaned up:

  • /opt/rtcds/caltech/c1/userapps/trunk/cds/c1/models/c1uct.mdl was linked in, but that model doesn't seem to exist anymore, so I removed the link.
  • a couple of things were linked from /opt/rtcds/caltech/c1/userapps/trunk instead of /opt/rtcds/caltech/c1/userapps/release, so I fixed those links.
  • /opt/rtcds/caltech/c1/userapps/release/cds/test/models/llo/l1isctest.mdl was not checked out, so I checked it out and fixed the link (this model should really be named something different if it is of common use, or we plan on using it at the 40m).

 

  4763   Mon May 23 18:16:42 2011 KojiConfigurationLSCPOY55 installed

The DC Transimpedance of POP55 was increased from 50 Ohm to 10010 Ohm. There is the offset of 46mV. This should be cancelled in the CDS.

Quote:

Last Saturday the POY55 RFPD (see this entry) was installed on the ITMY optical bench for the trial of the DRMI locking.

Since the amount of the light coming into the diode is tiny, the DC monitor showed ~ 3 mV even when the PRC was locked to the carrier.

In order to amplify the tiny RF signal from the photo diode a ZHL amplifier was installed next to the RFPD. The RF amp is sitting on delrin posts for insulation from the table.

 

  4762   Mon May 23 18:10:41 2011 kiwamuUpdateLSCf2p filters on PRM : not good

During the DRMI trial I noticed that the f2p filters on PRM is not quite effective (i.e. pushing PRM in POS direction makes misalignments).

I checked the f2p filters in an easy way. I pushed POS at 0.01 Hz with an amplitude of 1000 counts and looked at the oplev error signals with / without the f2p filters.

The picture below is a time series of the POS excitation, the oplev's PITCH and YAW error signals.

You can see there still is a big coupling from POS to YAW after the f2p filters were enabled. (Its supposed to be like this)

I will redo the f2p measurement on PRM.

f2p_PRM.png

  4761   Mon May 23 14:28:23 2011 kiwamuConfigurationLSCPOY55 installed

Last Saturday the POY55 RFPD (see this entry) was installed on the ITMY optical bench for the trial of the DRMI locking.

Since the amount of the light coming into the diode is tiny, the DC monitor showed ~ 3 mV even when the PRC was locked to the carrier.

In order to amplify the tiny RF signal from the photo diode a ZHL amplifier was installed next to the RFPD. The RF amp is sitting on delrin posts for insulation from the table.

POY55.png

  4760   Mon May 23 12:27:26 2011 kiwamuUpdateLSCDRMI trial : details

(PRMI locking with slightly misaligned SRM)

 First I tried locking PRC and MICH with a little bit misaligned SRM. This condition allowed me to search for a good signal port for SRC.

In this locking, REFL11_I was used to lock PRC and AS55_Q was used for MICH. This is the same scheme as the current PRMI locking.

Since the alignment of SRM was close to the good alignment, I expected to see fringes from SRC in some signal ports (i.e. REFL55, POY55 and so on).

Sometimes a fringe of SRC disturbed AS55_Q and broke the MICH locking, so I had to carefully misalign SRM so that the SRC fringes are small enough to maintain the lock of MICH.

 

(Looking for a good signal port for SRC)

 After I locked the PRMI with slightly misaligned SRM, I started looking for a good signal port for SRC.

At the beginning I tried finding a good SRC port by shaking SRM at 100 Hz and looked at the power spectra of all the available LSC signals.

I was expecting to see a 100 Hz peak in the spectra, but this technique didn't work well because SRC wasn't within the linear range and hence didn't produce linear signals.

So I didn't see any strong signals at 100 Hz and finally gave up this technique.

Then I started looking for a PDH-like signal in time series and immediately found AS55_I showed large PDH-like signals.

So I started using the AS55_I for the SRC locking and eventually succeeded.
 

 

(Two tips for the DRMI locking)

During the locking of DRMI, I found two tips that made the locking quite smooth.

 - Triggered locking

   Since every LSC signal ports showed large signals from PRC somehow, feeding back the signals made the suspensions crazy.

   So I used triggered locking for the PRC and MICH locking to avoid unwanted kicks on BS and PRM.

   If  the DC of REFL goes above a certain level, the control of  PRC starts. Also if the DC of AS goes below a certain level the control of MICH starts.

  These triggers make the lock smoother.

 - Do not use resonant gain filters

  This is really a stupid tip. When I was trying to lock MICH, the lock became quite difficult for some reasons.

  It looked there was an oscillation at 3 Hz every time the MICH control started. It turned out that a 3 Hz resonant gain filter had been making it difficult.

  All the resonant gain filters should be off when a lock acquisition is taken place.

Quote from #4759

Eventually the DRMI was locked.

More details will be reported in the morning.

 

  4759   Mon May 23 00:36:51 2011 kiwamuUpdateLSCDRMI trial : sucess

Eventually the DRMI was locked.

I was struggling to find a good signal port for SRC over the weekend and finally found AS55_I worked somehow. I used :

   REFL11_I --> PRC

   AS55_Q   --> MICH

   AS55_I    --> SRC

A configuration script was prepared such that someone can try this configuration by clicking a button on the C1IFO_CONFIGURE.adl screen.

I don't think this signal extraction scheme is the best, but now we can find better signal ports by shaking each DOF and looking at each signal port.

More details will be reported in the morning.

Quote:

I will try with POY55 that Koji prepared today.

 

  4758   Sat May 21 17:02:38 2011 KojiUpdateElectronicsAlberto's 11MHz was modified to POP55MHz

- Resonant at 55MHz. The transimpedance is 258Ohm. That is about half of REFL55 (don't know why).

- 11MHz&110MHz notch

- The 200MHz oscillation of MAX4106 was damped by the same recipe as REFL11.

POP55_transimpedance.pdf

 

Attachment 1: POP55_schematic_110520_KA.pdf
POP55_schematic_110520_KA.pdf POP55_schematic_110520_KA.pdf
Attachment 2: POP55_transimpedance.pdf
POP55_transimpedance.pdf
  4757   Sat May 21 06:19:46 2011 kiwamuUpdateLSCDRMI trial : no luck

I will try with POY55 that Koji prepared today.

  4756   Fri May 20 11:37:44 2011 KojiUpdateElectronicsREFL11 fixed (REFL11 not working)

- Found the inductor which shunts the positive input of MAX4107 was not touching the ground.
This left the positive input level undetermined at DC. This was why MAX had been saturated.
The PCB has a cut, so it was surprising once the circuit worked.

- Resoldered the inductor to the ground. This made the circuit responding to the intensity-modulated beam.

- But the resonances and the notches were totally off, and the 200MHz oscillation has resurrected.

- Attached 40Ohm+22pF network between the neg-input of MAX and the gnd. This solved the oscillation.

- Made the tuning and the characterizations. The PD is on Kiwamu's desk and ready to go.

More to come later

Quote:
I took REFL11 out from the AS table for a health check because it wasn't working properly.

The symptoms were :

   - a big offset of ~ -3 V on the RF output. No RF signals.

   - The DC output seemed to be okay. It's been sensitive to light.

I did a quick check and confirmed that +/- 5V were correctly supplied to the op-amps.

It looks that the last stage (MAX4107) is saturated for some reasons. Need more inspections.

At the moment the REFL11 RFPD is on the bench of the Jenne laser.

 

Attachment 1: REFL11_transimpedance.pdf
REFL11_transimpedance.pdf
  4755   Fri May 20 05:41:22 2011 SureshUpdateRF SystemDemod board measurements

The POP110 board which had the large Amplitude and Phase unbalances was examined today.  It turned out that there was some stray solder which had connected the Sum port of the PSCQ-2-120 splitter to its body (ground).  After I removed that the amplitude unbalance was 0.3dB however the phase was 105deg.  The phase reduced to 90 deg only if the power on the splitter is around 19 dBm.  So removed the AT1 (10dB attenuator) and the phase unbalance dropped to 91 deg.  However this is not a sustainable solution as the ERA-5 max output is about 19.5 dBm.

As this is a side band power monitor (and not a length sensing RFPD), we can make do with a poorer phase.  I will therefore replace the AT1 and adjust the residual phase with cable delay lines.  

Quote:

I measured the amplitude and phase imbalances of the demod boards which have been modified.  This is just a basic health check.  We hope to use the script that Kiwamu is developing for a more accurate test.  The script can also use these measurements as a sanity check.  POP110  requires some further attention. 

Demod_Board_measurements.png

 

The RF distribution box outputs corresponding to the demod board (eg. AS55_LO --> AS55_demod) were used as LO sources.  The RF signal was generated with a Marconi and held a kHz away from the LO frequency.  The amplitude and phase unbalance were measured with SR785.  The RF Power meter was used to check the LO power in each case.

 

 

 

  4754   Fri May 20 03:29:04 2011 kiwamuUpdateCDSBinary IO box on 1X5 : LEDs off

[Steve / Kiwamu]

 When Steve was working on the strain reliefs on 1X5 he found that some LEDs on the back side of the binary IO boxes were off.

There are 4 binary IO boxes and their power are directly supplied from Sorensens. According to the display of the Sorensens, the power are correctly generated.

Steve and I checked a picture of the boxes taken before he started working and we found it's been like this.

It might be just a problem of the LEDs or the fuses are blown, but anyway it needs an inspection.

Here is a picture of the back side of the boxes. You can see some LEDs are on and some are off.

DSC_3050_small.jpg

  4753   Fri May 20 03:01:17 2011 kiwamuUpdateLSClocking status

(PRMI locking)

Since REFL11 has gone I tried locking the PRMI with combination of REFL55 and AS55.

Without any pain the lock of PRMI was achieved successfully. AS55 was used to sense MICH and REFL55 was used for PRC.

 

(scripting)

Additionally I was modifying several scripts which are invoked from C1IFO_CONFIGURE.adl. Some details about the scripts will be uploaded on the wiki later.

An important thing is that now we are able to use the "restore" commands for the Y arm, X arm, Michelson and PRM locking.

The scripts will automatically acquire the lock of each DOF.  The image below is just a screen shot of the medm screen where you can call the scripts.

Screen_shot_2011-05-20_at_2.50.00.png

 

( Still to do)

   * PRM actuator response measurement

   * PRC noise budget

   * MICH-PRC actuator decoupling

  4752   Fri May 20 01:02:50 2011 loose connection hunterUpdateSUSloose connection on ETMY rack

The UL signal of the shadow sensor on ETMY went to zero this evening.

This was due to a loose connection on the cross connection board on the 1Y4 rack.

In order to make them tighten, a combination of stand-offs and screws were installed on the connectors. They won't be loose any more.

ETMY_loose.png

  4751   Thu May 19 19:41:17 2011 SureshUpdateRF SystemDAQ channel Assignments for RF PDs

[Kiwamu, Suresh]

In trying to keep the wiring of the LSC rack as neat as possible, we came up with the following channel assignments of the RF PD signals. 

PDI = PD Interface, The PD Interface D-type connectors (#1 to #12) are numbered:  Top -> Bottom and Left -> Right in ascending order. 

The Analog channels on the LSC Whitening Filter boards are numbered similarly, 1 to 32 in four sets.

1Y2_Rack_Layout.png

  4750   Thu May 19 17:53:03 2011 steveUpdateCDSAA filter box modified at 1X5

Quote:

[Steve, Joe]

Steve pulled the top AA filter box from 1X5 which handled some of the suspensions channels.  We turned off all the watchdogs before pulling it out, as well as recorded which cables were connected to which inputs.

The case  is undergoing a structural modification to have the ADC adapter card which previously was loosely connected via cables, securely attached to the case.

Steve still wants to do some cabling in the rack while the box is out, and will return it this afternoon once he has finished that.

 Job is done. Sus damping are back on. Cabling-strain reliefing are  not finished yet at 1X5 and 1X4

Attachment 1: P1070727.JPG
P1070727.JPG
  4749   Thu May 19 16:46:20 2011 kiwamuUpdateLSCc1lsc model : input channels rearanged

According to Suresh's LSC rack design I rearranged the input channels of the c1lsc model such that the analog signals and the ADC channels are nicely matched.

Also I updated the c1lsc model in the svn with a help from Joe. The picture below is a screen shot of the input channels in the model file after I edited it.

c1lsc.png

  4748   Thu May 19 12:09:41 2011 josephbUpdateCDSAA filter box pulled from 1X5, optic suspensions currently off

[Steve, Joe]

Steve pulled the top AA filter box from 1X5 which handled some of the suspensions channels.  We turned off all the watchdogs before pulling it out, as well as recorded which cables were connected to which inputs.

The case  is undergoing a structural modification to have the ADC adapter card which previously was loosely connected via cables, securely attached to the case.

Steve still wants to do some cabling in the rack while the box is out, and will return it this afternoon once he has finished that.

  4747   Thu May 19 03:13:54 2011 kiwamuUpdateElectronicsREFL11 not working

I took REFL11 out from the AS table for a health check because it wasn't working properly.

The symptoms were :

   - a big offset of ~ -3 V on the RF output. No RF signals.

   - The DC output seemed to be okay. It's been sensitive to light.

I did a quick check and confirmed that +/- 5V were correctly supplied to the op-amps.

It looks that the last stage (MAX4107) is saturated for some reasons. Need more inspections.

At the moment the REFL11 RFPD is on the bench of the Jenne laser.

 

  4746   Thu May 19 00:23:44 2011 ranaUpdateCDSdiagonalization of MC input matrix

 I've moved all of my SOS peak fitting stuff into the scripts area so that Leo can make it better:

/cvs/cds/rtcds/caltech/c1/scripts/SUS/peakFit

findPeaks.m gets the data and makes the fitted spectra that I put in the previous entry.

findMatrix.m is the barely started script that ought to take the TF data and output the matrix to the MEDM screen.

  4745   Thu May 19 00:22:21 2011 ranaUpdateelogRestarted, Italian-style

Quote:

Aka, from a hotel in Pisa.

 Restarted Thu May 19 00:21:49 2011 to recover from Jenne's Italian terrorism.

  4744   Thu May 19 00:15:20 2011 JenneUpdateelogRestarted, Italian-style

Aka, from a hotel in Pisa.

  4743   Wed May 18 22:11:31 2011 ranaConfigurationLSCNew digital lockin added to LSC model/screen

Untitled.png

 This is OK....but, the input matrix should come from the same place as the regular input matrix: i.e. it should be just another row like CARM, DARM, etc. rather than have its own screen.

Also, I think a nice mod to all the matrices would be if the ORANGE triangle was only visible when there's a signal flowing through it.

  4742   Wed May 18 18:48:46 2011 kiwamuUpdateElectronicsidentification of RFPD interface cables

[Haixing / Kiwamu]

 As a part of the Wednesday's cabling work, we spent some times for identifying the RFPD interface cables.

The RFPD interface cables are made of a 15 pin flat cable, containing DC power conductors for the RFPDs and the DC signal path.

The list below is the status of the interface cables.

 

- - - - RFPD name, (cable status) - - - -

- REFL11 (identified and labeling done)

- REFL33 (identified and labeling done)

- REFL55 (identified and labeling done)

- REFL165 (no cable found)

- AS55 (identified and labeling done)

- AS165 (identified and labeling done)

- POP22/110 (identified and labeling done)

- POX11 (identified and labeling done)

- POY11 (identified and labeling done)

- POY55 (identified and labeling done)

We still have two cables which are not yet identified. Their heads are around the LSC rack and labeled 'unidentified'

  4741   Wed May 18 18:33:46 2011 AidanUpdateelogrestarted elog with script
  4740   Wed May 18 17:06:39 2011 SureshUpdateRF SystemDC power supplies for the RF generation box in place

 

I have checked the voltages on the connector.  They are okay and I have plugged in the Sorensen power into the RF Source.  The ground reference for the Sorensens comes from the 1X2 Rack ground reference lines on the south side of the rack. 

I looked for the OMC ground reference. Could not find one on either of the the OMC half racks.

Quote:

[Koji, Steve]

DC power supplies for the RF generation box are now in place. They are the top two of the 6 Sorensens in the OMC short rack next to 1X2.

We made the connections as we did for the RF distribution box, the power supplies labele, and the cables strain-relieved.

The power supply is not yet connected to the actual RF generation box. This should be done by Suresh or someone with the supervision of him.

Note:
We have two +18V supply on the short OMC rack, in total. One is for the RF source, the other is for the OMC PZTs, whitening, etc.
This is to avoid unnecessary ground loop
although the grounding situation of the OMC side is not known to me.

 

  4739   Wed May 18 16:52:23 2011 SureshUpdateRF SystemCables for AS11 PD are in place

[Larisa, Suresh]

All the cables needed for the AS11 PD are in place... the heliax cable runs from the AS table to the PSL rack.  The LO and RF cables to demod board as well as the I and Q cables into the LSC Whitening board are connected.

The cables get rather densely packed when the LSC Whitening filter sits between the PD Interface Board and the LSC AA filter board.  This makes it difficult to access the SMA connectors on the LSC whitening filter.  So we shifted the LSC Whitening and AA Filter boards one slot to the right.  The LSC rack looks like this just now.  We have also shifted the binary cables at the back of the Eurocart by one slot so the same cables are associated with the cards.

1Y2_Rack_Layout_short-term.png

 

 

  4738   Wed May 18 15:54:50 2011 KojiUpdateRF SystemDC power supplies for the RF generation box in place

[Koji, Steve]

DC power supplies for the RF generation box are now in place. They are the top two of the 6 Sorensens in the OMC short rack next to 1X2.

We made the connections as we did for the RF distribution box, the power supplies labele, and the cables strain-relieved.

The power supply is not yet connected to the actual RF generation box. This should be done by Suresh or someone with the supervision of him.

Note:
We have two +18V supply on the short OMC rack, in total. One is for the RF source, the other is for the OMC PZTs, whitening, etc.
This is to avoid unnecessary ground loop
although the grounding situation of the OMC side is not known to me.

  4737   Wed May 18 07:18:15 2011 SureshUpdateRF SystemPOY55_Demod board Hardware change completed

The ~350 Hz noted in the elog below was traced to an RF modulation of the 11 MHz sideband.  This modulation was set up in the Marconi which is currently supplying the 11 MHz local oscillator signal to the RF source.  lt was used during the MC length study completed last week by Valera and Ryan.  The frequency measured was 322 Hz.

As we do not require this any longer, I have switched off this modulation.

 

 

Quote:

The Demod board with S. No. 022 (being used earlier as REFL11) has been modified.  It now has SCLF-65 as its input LP filter on the PD input line and a PQW-2-90 power splitter.  The unit functioning okay (I and Q signals are 90 deg apart.

The loss of Q output was traced to a possible loose solder joint and we now have both the I and Q signals after resoldering all components in the vicinity of U7 (Ref Schematic of D990511)

There is a strong oscillation around 350Hz present on I and Q signals of both REFL55_Demod and POY55_Demod.  Don't know the source. 

We have run out of power splitters to continue with the Demod board modification. We do not currently have an AS11_Demod board.  All the others are in place and ready for the I<->Q phase angle measurement.

In summary we now have the following Demod boards in place:

[ REFL11, POY11, REFL55, AS55, POY55, POY22, POY110]_Demod

 

 

  4736   Wed May 18 07:13:00 2011 SureshUpdateRF SystemDemod board measurements

I measured the amplitude and phase imbalances of the demod boards which have been modified.  This is just a basic health check.  We hope to use the script that Kiwamu is developing for a more accurate test.  The script can also use these measurements as a sanity check.  POP110  requires some further attention. 

Demod_Board_measurements.png

 

The RF distribution box outputs corresponding to the demod board (eg. AS55_LO --> AS55_demod) were used as LO sources.  The RF signal was generated with a Marconi and held a kHz away from the LO frequency.  The amplitude and phase unbalance were measured with SR785.  The RF Power meter was used to check the LO power in each case.

 

 

Attachment 1: Demod_Board_measurements.png
Demod_Board_measurements.png
  4735   Tue May 17 22:50:00 2011 JamieConfigurationLSCNew digital lockin added to LSC model/screen

I added a lockin to the LSC model, and added it's corresponding control to the LSC screen.  Here's is what I added to the LSC model:

lsc-mdl-lockin.png

On the left are the froms from the normalized RF PD outputs. Those go into a matrix, whose single row goes into the lockin signal input. The clock output from the lockin goes into the LSC output matrix (last row).

Here is what I added to the LSC master screen:

lsc-screen-lockin.png

I also modified the lockin overview screen:

lsc-lockin-screen.png

I think this new screen looks a lot cleaner.  Maybe we could start using this one as a template for lockin screens.

  4734   Tue May 17 19:38:32 2011 kiwamuUpdateSUSwrong connection on 1X5

Today Steve was working around the 1X5 rack to strain relief the cable jungles and the jungle is now getting less jungle.

During the work he disconnected and reconnected some cables.

So for a doublecheck I checked all the suspensions to see if the suspensions are still healthy or not.

Aha, then I found a mistake.

 

See the pictures below. It's a very subtle difference. This wrong connection prevented MC1 and MC3 from damping.

wrong.png correct.png

  4733   Tue May 17 18:09:13 2011 Jamie, KiwamuConfigurationCDSc1sus and c1auxey crashed, rebooted

c1sus and c1auxey crashed, required hard reboot

For some reason, we found that c1sus and c1auxey were completely unresponsive.  We went out and gave them a hard reset, which brought them back up with no problems.

This appears to be related to a very similar problem report by Kiwamu just a couple of days ago, where c1lsc crashed after editing the C1LSC.ini and restarting the daqd process, which is exactly what I just did (see my previous log).  What could be causing this?

  4732   Tue May 17 17:01:22 2011 jamieConfigurationCDSUpdate LSC channels from _DAQ to _DQ

As of RCG version 2.1, recorded channels use the suffix "_DQ", instead of "_DAQ".  I just rebuilt and installed the c1lsc model, which changed the channel names, therefore hosing the old daq channel ini file.  Here's what I did, and how I fixed it:

$ ssh c1lsc
$ cd /opt/rtcds/caltech/c1/core/trunk
$ make c1lsc
$ make install-c1lsc
$ cd /opt/rtcds/caltech/c1/scripts
$ ./startc1lsc
$ cd /opt/rtcds/caltech/c1/chans/daq
$ cat archive/C1LSC_110517_152411.ini | sed "s/_DAQ/_DQ/g" >C1LSC.ini
$ telnet fb 8087
daqd> shutdown

 

  4731   Tue May 17 16:53:22 2011 kiwamuUpdateSUSoplev sign was wrong on ETMY

Last night I found that the sign of the oplev control of PITCH on ETMY was wrong. I flipped it to the correct sign.

We've been locking the Y arm by feeding a signal back to ITMY  because pushing ETMY somehow made the lock unstable in the angular motion.

After the correction of the oplev contol sign, I was able to keep the lock robustly by pushing ETMY.

  4730   Tue May 17 11:45:20 2011 JamieConfigurationCDSpurged non-c1 site files from rtcds checkout of cds_user_apps
I purged all of the working copy checkouts of site files for all sites that are not c1 from the rtcds cds_user_apps working directory (/opt/rtcds/caltech/c1/userapps/trunk). I first checked that there were no outstanding changes, and then did the following (in bash):
cd /opt/rtcds/caltech/c1/userapps
rm -rf trunk
svn update --depth=files trunk
svn update --depth=empty trunk/{CDS,ISI,ISC,PSL,SUS}
svn update trunk/{CDS,ISI,ISC,PSL,SUS}/{c1,common}
  4729   Tue May 17 01:05:56 2011 kiwamuUpdateLSCAsymmetry measurement prep : recentering works

I re-centered beams on several PDs and a camera including :

  AS55, ETMY_QPD, TRY and ETMYT_CCD.

 

The most important one was AS55.

When I was locking each arm I found that the error signal from AS55 was very coupled to the angular motion of the arms.

I checked the beam on the AS55 RFPD and found the beam on the edge of the photo diode. This is possibly because Valera and I had been touching the input beam alignment.

At that time the DC signal from AS55 without aligning PRM and SRM was about 5 mV.

Adjusting the beam position by a steering mirror brought the DC signal up to 20 mV.

Then the lock of each arm became more stable.

  4728   Mon May 16 17:11:28 2011 steveUpdateElectronicsstrain relieved 1X1

1X1 is strain relieved in the back. I will use similar approach on the rest of the racks .

Attachment 1: P1070724.JPG
P1070724.JPG
  4727   Mon May 16 15:46:05 2011 KojiUpdateGeneraloptical layout viewable on the online version of autocad

I have made my own AutoCAD WS account and put the latest 40m layout.
AutoCAD WS is a free cloud service which enable us to browse/edit the DWG files.

You can view/play/print it from the following link even without making the account.

https://www.autocadws.com/main/publish?link=RVFmNEtOd3EzNVFtRXcx

If you make your account you can actually save it although the editing capability is somewhat limited.

Notes:
- It needs the Flash plug-in and Internet connection.
- I think Safari is the most stable platform on Mac. (i.e. I had some malfunctions with Firefox and Chrome)

  4726   Mon May 16 11:47:59 2011 kiwamuUpdateASSc1ass update part II

The medm screen for c1ass started being modified to be more user-friendly.

The modification is still ongoing, but the goal is to make a screen which anyone can easily understand and play with.

 

Still to do : ( need a volunteer )

- Modification of the screens

- Commission the input beam and X-arm servos

- Make scripts for X-arm

- Measure the PZT mirrors' matrix for the translation and angle

c1ass.png

Quote from #4709

Here the status of the dither alignment or c1ass:

Still to do:

- Commission the input beam and X-arm servos

- Make scripts for X-arm

 

ELOG V3.1.3-